OpenStudy (anonymous):

Write a formula for this sequence: 19, 11, 3, -5,...

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

It would be x-8...

OpenStudy (anonymous):

witch one? A. 27 - 8n B. -8n + 19 C. -5 + 8n

OpenStudy (anonymous):

If n starts at 1 and increases, then the answer is A. If n starts at 0 and increases, then the answer is B. I believe the answer you are looking for is A, since most sequences go from n=1 to n=inf

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Find the 7th term in this sequence: tn=8.5n-30 A. -21.5 B. 29.5 C. 4 D. 38

OpenStudy (anonymous):

Since you want the 7th term, you should substitute 7 into n, and that would give you 8.5(7)-30=59.5-30=29.5. Your answer would be B. 29.5.

OpenStudy (anonymous):

The easiest way to approach this problem is to note that the you can generally put the term number into the equation. In this case, 8.5*7-30=29.5.

OpenStudy (anonymous):

thanks sooo much guys!

OpenStudy (anonymous):

Write a formula for this sequence. 50, 83, 116, 149,... A. 33n + 50 B. 33n + 17 C. 149 - 33n

OpenStudy (anonymous):

So, begin with the first answer. Note that for A, if you put in term n=1, you get 83. This is not the first term, so A is wrong. For B, if you put in term n=1, you get 50. That is the first term, so B could be correct. Now note that each term increases by 33, so B is probably the right answer. Now check C. If you put in n=1, you get 116, which is not the first term. So B is the answer.

OpenStudy (anonymous):

omg u explained it way better then the review did :) thx!

OpenStudy (anonymous):

Find the slope for the table of values below. x 1 2 3 4 y 8.5 11 13.5 16 A. 2.5 B. 4 C. 1 D. 8.5

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

So the easiest way to find the slope is to know that rise/run is equal to the slope. Since the change in y is the rise and the change in x is the run, we can determine the slope from the table. The change in y is 2.5, because the numbers increase by 2.5 every time x increases by 1. The change in x is 1. So, the slope would be change in y/change in x, which is 2.5/1 or 2.5. A

OpenStudy (anonymous):

thankyou!

OpenStudy (anonymous):

Find the equation for the table. x 1 2 3 4 y 8.5 11 13.5 16 this would be the same rule right??

OpenStudy (anonymous):

Well now you know the slope. If you are familiar with the general equation y=mx+b, you can see that m stands for slope. We solved for slope in the previous example, so we just need b which is the y intercept.

OpenStudy (anonymous):

Yes, this would be the same rule for finding the slope.

OpenStudy (anonymous):

ok i understand that one.... but i dont get this one Predict the y-value for x = 6. x 1 2 3 4 y 8.5 11 13.5 16 A. 19.5 B. 21 C. 23.5 D. 26

OpenStudy (anonymous):

Ok so to predict the y-value, you just have to extrapolate. Note that every time x increases by 1, y increases by 2.5. So if you keep adding 2.5 to y and extend the table, you'll reach x=6.

OpenStudy (anonymous):

so what would the answer be?

OpenStudy (anonymous):

So if you add 1 to x to make it x=5, then you have to add 2.5 to y to make it y=18.5. Now add 1 to x to make it x=6, and add 2.5 to y to make it y=21. So when x=6, y=21.

OpenStudy (anonymous):

ooooooooh

OpenStudy (anonymous):

Is the following sequence arithmetic? 11, 15, 19, 23,... A. yes B. no